OpenStudy (anonymous):

if i travel 1,126.4 inches per second, how fast do i travel in miles per hour

OpenStudy (anonymous):

how many inches are there in one mile? how many seconds are there in one hour?

OpenStudy (anonymous):

60 seconds in one hour and 63360 inches per mile

OpenStudy (anonymous):

okay just follow the unit cancellation method (that's how i call it lol)

OpenStudy (anonymous):

give me a sec to type it

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

\(\sf \Large (\frac{1,126.4\ inches}{second}) ( \frac {1 mile}{63360 inches})( \frac {60 sec}{1 hour})\).. cancel inches, cancel seconds.. you'll end up with \(\sf 1,126.4 \times 60 \div 63360= [ans] \ mile/hour\)

OpenStudy (anonymous):

so about 1.067mph ?

OpenStudy (anonymous):

yep